9(b) Consider the system of equations (5)

11 2x1 + x2 + 4x3 =

3x1 + x2 + 5x3 =14

A feasible solution is x1 = ,2 x2 = ,3 x3 = .1 Reduce this feasible solution to a basic

feasible solution.


1
Expert's answer
2021-05-10T13:40:41-0400

"2x_1+x_2=11-4x_3"

"3x_1+x_2=14-5x_3"

"\Delta=\begin{vmatrix}
 2 & 1 \\
 3 & 1
\end{vmatrix}=-1"

"\Delta_1=\begin{vmatrix}
 11-4x_3 & 1 \\
 14-5x_3 & 1
\end{vmatrix}=x_3-3"

"\Delta_2=\begin{vmatrix}
 2 & 11-4x_3 \\
 3 & 14-5x_3
\end{vmatrix}=2x_3-5"

"x_1=\Delta_1/\Delta=3-x_3"

"x_2=\Delta_2/\Delta=5-2x_3"